OpenStudy (anonymous):

can someone explain to me the process how to Factor? example factor 27y^3+x^9. and Factor by grouping: xw+3x−6w−18

OpenStudy (amistre64):

sum of cubes ..... seems to be the first one. and ive never got the hang of grouping :/

OpenStudy (amistre64):

\[a^3+b^3=(a+b)(a^2-ab+b^2)\] this is just one of those rules that you have to memorize

OpenStudy (amistre64):

xw+3x−6w−18; this is usually accomplished by forming an addition of 2 groups: (xw+3x) + (−6w−18), take out common factors from each group ^^ ^^ x -6 x (w+3) -6 (w+3) ; with any luck, this produces another common factor of the 2 groups (w+3) (x-6)

OpenStudy (anonymous):

I got it up to here x(w+3) -6(w+3) , but how do you get (w+3) (x-6)?

OpenStudy (amistre64):

what is the common factor of: x(w+3) -6(w+3) we can clean this up by letting a = w+3 xa -6a, do you agree that combining like terms of a, we get: (x-6) a ??

OpenStudy (amistre64):

if so, then we can just replace a with what we defined it to be to begin with (x-6) a = (x-6) (w+3)

OpenStudy (anonymous):

OK i get it, the common factor for xa-6a is a so a(x-6)

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

amistre64 could you help i got lost in factoring 27y^3+x^9. i used a3+b3=(a+b)(a2−ab+b2) but i'm not getting anywhere

OpenStudy (amistre64):

a = 3y, a^3 = 27y^3 b = x^3, b^3 = x^9 thats all that amounts to

OpenStudy (amistre64):

recall that when we raise a product to an exponent, we raise all the parts as well. \[(kx)^3=kn*kn*kn=k^3n^3\]

OpenStudy (amistre64):

lol, umm either change x to n, or n to x i seem to have been thinking of 2 different ideas at ones

OpenStudy (amistre64):

\[(kx)^n=kx*kx*kx*...*kx:n~times=k^nx^n\] \[(kx)^n=k^nx^n\]

OpenStudy (anonymous):

ok thanks